# Finite Measures of Maximal Entropy for an Open Set of Partially Hyperbolic Diffeomorphisms

### v1

@@ 2.4.2. Unstable topological entropy and the variational principle
$h_{\text{top }}^{u}(f)=\lim_{\delta\rightarrow 0}\sup_{x\in M}h_{\text{top }}^{u}\left(f,\overline{W^{u}(x,\delta)}\right),$
@@
$h_{\text{top }}^{u}\left(f,\overline{W^{u}(x,\delta)}\right)=\lim_{\epsilon\rightarrow 0}\limsup_{n\rightarrow\infty}\frac{1}{n}\log N^{u}(f,\epsilon,n,x,\delta).$
Analogously, if the dimension of the stable bundle $E^{s}_{f}$ is greater than or equal to 1, we can define the stable entropy for any $\mu\in{\mathcal{M}}_{f}$ as $h^{s}(f):=h^{u}_{\mu}(f^{-1})$.
Similar to the traditional definition of entropy, we can establish a connection between metric unstable entropy and unstable topological entropy through a variational principle. Specifically, according to [10, Theorem D], for a $C^{1}$-partially hyperbolic diffeomorphism $f:M\rightarrow M$, the following relation holds:
$h^{u}_{top}(f)=\sup\{h^{u}_{\mu}(f):\mu\in{\mathcal{M}}_{f}\}\quad\mbox{and}\quad h^{u}_{top}(f)=\sup\{h^{u}_{\nu}(f):\nu\in{\mathcal{M}}_{f}^{e}\}.$
Another way to define topological unstable entropy is by considering the unstable volume growth given by Hua, Saghin, and Xia ([12]), reminiscent of the works by Yomdin and Newhouse ([34, 20]). In [10, Theorem C] it is shown that the unstable topological entropy as defined here coincides with the unstable volume growth.
@@ 2.5. Hausdorff distance
In this section we will define the Hausdorff distance, we follow [1, Section 7.3].
@@
Let $(K_{n}){n\in{\mathbb{N}}}$ be a convergent sequence of compact subsets of the compact metric space $(M,d)$ converging to $K$ in the Hausdorff Metric. If $x_{n}\in K_{n}$ forms a convergent sequence in $(M,d)$ converging to $x$, then $x\in K$. Moreover, if $(\mu_{k})_{k\in{\mathbb{N}}}$ are probability measures converging to $\mu$ in the weak * topology, then $\mu(K)\geq\limsup(\mu_{n}(K_{n}))$.
@@ 3. Consequences of $h^{u}(f)-h^{s}(f)>0$
Consider $f:M\to M$ such that $h^{u}(f)>h^{s}(f)$. If $\mu$ is an ergodic measure such that $h_{\mu}(f)>h^{s}(f)$, then $\lambda_{c}(f,\mu)<h^{s}(f)-h_{\mu}(f)<0$.
@@
Suppose that the central Lyapunov exponent of $\mu$ is non-negative, i.e., $\lambda^{c}(\mu,f)\geq 0$. Applying (3.2) to the stable metric entropy, we have
$h_{\mu}(f)=h^{s}_{\mu}(f)$
and by hypothesis, $h_{\mu}(f)>h^{s}(f)$, therefore $h_{\mu}^{s}(f)>h^{s}(f)$, which contradicts the variational principle for stable entropy. ∎
By $h_{\mu}(f)>h^{s}(f)$ and Claim 1, we have that $\lambda_{c}(f,\mu)<0$, and therefore by 3.1,
@@
It is worth noting that under the hypotheses of Lemma 3.1, every ergodic measure with large entropy is hyperbolic and the metric entropy coincides with the unstable metric entropy. Furthermore, the central Lyapunov exponent is negative, so its $s$-index is $\operatorname{dim}(E^{s})+1$. This leads us to inquire whether two measures with large entropy exhibit homoclinic relation as defined in the definition 2.5.
@@
Observe that the condition $P^{u}(f,\phi)>P^{s}(f,\phi)$ is not sufficient to guarantee positive entropy for the equilibrium states of $(f,\phi)$. When $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$, every measure of maximal entropy $\mu$ satisfies $P_{\mu}(f,\phi)>h^{s}(f)+\sup\phi$. So,
@@
Let $\phi:M\to{\mathbb{R}}$ be a continuous potential such that $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$. If $\mu$ is an ergodic measure such that $P_{\mu}(f,\phi)>h^{s}(f)+\sup\phi$ then $\lambda_{c}(f,\mu)<h_{\mu}^{s}(f)-h^{s}(f)<0$.
@@
If $h^{u}(f)>a_{1}>a_{2}>h^{s}(f)$, then there exists a $C^{1}$ neighborhood ${{\mathcal{U}}}$ of $f$ such that for every $g\in{{\mathcal{U}}}\cap{{\rm Diff}}^{1+}(M)$, $h^{u}(g)>a_{1}>a_{2}>h^{s}(f)$, the function $h:{{\rm Diff}}^{1+}(M)\cap\mathcal{U}\to\mathbb{R}$ is continuous.
@@
The lemma is a consequence of [18, Corollary 3.6] and [18, Corollary 3.7]. ∎
@@ 4. Invariant manifolds for measures with large entropy
In this section, we will investigate the measurement of the set of points with invariant manifolds of uniform size from the perspective of measures with high entropy.
@@ 4.1. Invariant Manifolds with Uniform Size
We see that ergodic measures with large entropy are hyperbolic measures, see Lemma 3.1. Almost every point of a hyperbolic measure possesses stable and unstable invariant manifolds, see Section 2.3.2. However, a crucial point to note is the absence of uniform control over the size of these invariant manifolds. Consequently, there may exist points that are arbitrarily close to each other such that their corresponding unstable and stable manifolds may not intersect. This suggests the potential existence of an infinite number of maximal entropy measures that are not homoclinically related. To avoid this kind of problem, in this section, we construct, for partially hyperbolic diffeomorphisms with central dimension one a set where the Pesin invariant manifolds have uniform size.
Recall that $f:M\to M$ denotes a $C^{1+}$ partially hyperbolic diffeomorphism with $1$-dimensional central direction with non trivial stable and unstable bundles, that is, ${\operatorname{dim}}(E^{\sigma})>0(\sigma=s,u)$. Given $01>\rho\geq\xi>$ 0 where $\xi$ is as in 2.10, define
(4.1) $\Lambda_{\rho}(f)=\Lambda_{\rho}:=\{x\in M:\forall n\geq 1\mbox{ it holds }\|Df^{n}(x)_{|E^{c}}\|\leq\rho^{n}\}$
@@
and the splitting $E^{cs}(x)\oplus E^{u}(x)$ varies continuously. Therefore, the classical result presented in [11] guarantees the existence of invariant Pesin manifolds of uniform size. More precisely, there exists $\delta_{r}=\delta_{r}(\rho)$ and $\delta_{l}=\delta_{l}(\rho)$ such that every $x\in\Lambda_{\rho}(f)$ unstable and stable manifold have a size larger than $\delta_{r}$ and if $x,y\in\Lambda_{\rho}(f)$ and $d(x,y)<\delta_{l}$ then $W^{u/s}_{\delta_{r}}(x)\pitchfork W^{s/u}_{\delta_{r}}(y)$ is non empty. Furthermore, the invariant manifolds exhibit continuous variations concerning both the diffeomorphisms involved and the chosen points. In other words, given $f$ and $g$ are $C^{1+}$ close enough in the $C^{1}$ topology, for any $x$ in $\Lambda_{\rho}(f)$ and $y$ in $\Lambda_{\rho}(g)$ where $x$ and $y$ are close, the invariant manifolds $W^{s/u}_{\delta_{r}}(x)$ are also close to $W^{s/u}_{\delta_{r}}(y)$ within the $C^{1}$ topology.
4.2. The measure of $\Lambda_{\rho}$
Now, we know that points in $\Lambda_{\rho}(f)$ have invariant manifolds of uniform size. We will use this set to prove that there exists a finite number of homoclinic classes containing measures with large entropy. For this, it is crucial that this set is relevant for measures with large entropy. Therefore, we aim to demonstrate that the set $\Lambda_{\rho}(f)$ possesses a positive measure for measures with significant entropy. To achieve our goals, we require the following version of the Pliss Lemma.
@@ 4.2. The measure of $\Lambda_{\rho}$
[6, Lemma 3.1] For any $\varepsilon>0$, $\alpha_{1}<\alpha_{2}$ and any sequence $(a_{i})\in(\alpha_{1},+\infty)^{{\mathbb{N}}}$ satisfying
@@
    for any $l\geq 1$ and $n>n_{l}$, one has $\frac{a_{n_{l}}+\cdots+a_{n-1}}{n-n_{l}}\leq\alpha_{2}+\varepsilon$
@@
We define $m(f):=\min_{x\in M}\{\log\|Df(x)_{|E^{c}_{x}}\|\}$. Note that $m(f)$ varies continuously on $f$ in the $C^{1}$ topology.
@@
Suppose that $h^{u}(f)>h^{s}(f)$. If $\mu\in{\mathcal{M}}_{f}^{e}$ with $h_{\mu}(f)>h^{s}(f)$ then for every $e^{\lambda_{c}(f,\mu)}<\rho<1$ we have that $\mu(\Lambda_{\rho})\geq\frac{\log(\rho)-\lambda_{c}(f,\mu)}{\log(\rho)-m(f)}$
@@
[…]
$\frac{1}{|n|}\sum_{j=0}^{n-1}\delta_{f^{-j}(x)}\to\mu,$
where $\delta_{x}$ is the Dirac measure at point $x$.
Since $h^{u}(f)>h^{s}(f)$ and $h_{\mu}(f)>h^{s}(f)$, Lemma 3.1 implies that $\lambda_{c}(f,\mu)<h^{s}(f)-h_{\mu}(f)<0$. Now, let $10<e^{\lambda_{c}(f,\mu)}<\rho<$ 1 and define the set
$\Lambda(f)=\Lambda_{\rho}\cap\Lambda_{\mu}\cap\mathcal{R}.$
Note that $\mu(\Lambda(f))=\mu(\Lambda_{\rho})$.
@@
Consequently, we obtain $\|Df^{n}(f^{n_{l}}(x))_{|E^{c}(f^{n_{l}}(x))}\|<\rho^{n}$ for every $n\in{\mathbb{N}}$. Thus, for each $l$, $f^{n_{l}}(x)\in\Lambda_{\mu}$. Applying Birkhoff’s theorem and the Pliss Lemma, we arrive at the following estimate:
(4.2) $\mu(\Lambda_{f})\geq\limsup_{l\to\infty}\frac{l}{n_{l}}\geq\frac{\log(\rho)-\lambda_{c}(f,\mu)}{\log(\rho)-m(f)}>0.$
@@
Let $\phi:M\to{\mathbb{R}}$ be a potential continuous such that $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$. For every $a$ such that $h^{u}(f)+\inf\phi>a>h^{s}(f)+\sup\phi$ and $\frac{h^{s}(f)-a+\inf\phi}{2}<\log(\rho)<1$ it holds
@@
Now, consider $\log(\rho)>\frac{h^{s}(f)+\sup\phi-a}{2}$. So, $e^{\lambda_{c}(g,\mu)}<\rho<1$, and therefore by Corollary 4.4:
Corollary 4.7.
Let $\phi:M\to\mathbb{R}$ be a potential continuous such that $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$. For every $a$ such that $h^{u}(f)+\inf\phi>a>h^{s}(f)+\sup\phi$ and $\frac{h^{s}(f)-a+\inf\phi}{2}<\log(\rho)<1$ it holds
  1. (a)
    For every $g\in{\mathcal{V}}$,
[…]
  2. (b)
Proof.
(4.8) $\lambda_{c}(g,\mu)\leq h^{u}(g)-h_{\mu}(g).$
(4.9) $h^{s}(f)<\frac{h^{s}(f)+a-\sup\phi}{2}.$
[…]
(4.10) $h_{\mu}(f)\geq a-\sup\phi.$
Applying 4.9 and 4.10 in 4.8 we get:
$\lambda_{c}(g,\mu)<\frac{h^{s}(f)+\sup\phi-a}{2}.$
Now, consider $\log(\rho)>\frac{h^{s}(f)+\sup\phi-a}{2}$. So, $e^{\lambda_{c}(g,\mu)}<\rho<1$, and therefore by Corollary 4.4:

### v3

@@ 2.4.2. Unstable topological entropy and the variational principle
$h^{u}(f)=\lim_{\delta\rightarrow 0}\sup_{x\in M}h^{u}\left(f,\overline{W^{u}(x,\delta)}\right),$
@@
$h^{u}\left(f,\overline{W^{u}(x,\delta)}\right)=\lim_{\epsilon\rightarrow 0}\limsup_{n\rightarrow\infty}\frac{1}{n}\log N^{u}(f,\epsilon,n,x,\delta).$
Analogously, if the dimension of the stable bundle $E^{s}_{f}$ is greater than or equal to 1, we can define the stable entropy for any $\mu\in{\mathcal{M}}_{f}$ as $h^{s}(f):=h^{u}(f^{-1})$.
Similar to the traditional definition of entropy, we can establish a connection between metric unstable entropy and unstable topological entropy through a variational principle. Specifically, according to [17, Theorem D], for a $C^{1}$-partially hyperbolic diffeomorphism $f:M\rightarrow M$, the following relation holds:
$h^{u}(f)=\sup\{h^{u}_{\mu}(f):\mu\in{\mathcal{M}}_{f}\}\quad\mbox{and}\quad h^{u}(f)=\sup\{h^{u}_{\nu}(f):\nu\in{\mathcal{M}}_{f}^{e}\}.$
An ergodic $f$ invariant measure satisfying $h_{\mu}(f)=h^{u}(f)$ is called an ergodic measure of maximal $u$-entropy.
Another way to define topological unstable entropy is by considering the unstable volume growth given by Hua, Saghin, and Xia ([19]), reminiscent of the works by Yomdin and Newhouse ([54, 32]). In [17, Theorem C] it is shown that the unstable topological entropy as defined here coincides with the unstable volume growth.
@@ 2.5. Hausdorff distance
In this section we will define the Hausdorff distance, we follow [2, Section 7.3].
@@
Let $(K_{n})_{n\in{\mathbb{N}}}$ be a convergent sequence of compact subsets of the compact metric space $(M,d)$ converging to $K$ in the Hausdorff Metric. If $x_{n}\in K_{n}$ forms a convergent sequence in $(M,d)$ converging to $x$, then $x\in K$. Moreover, if $(\mu_{k})_{k\in{\mathbb{N}}}$ are probability measures converging to $\mu$ in the weak * topology, then $\mu(K)\geq\limsup(\mu_{n}(K_{n}))$.
@@ 3. Consequences of $h^{u}(f)-h^{s}(f)>0$
Consider $f:M\to M$ such that $h^{u}(f)>h^{s}(f)$. If $\mu$ is an ergodic measure such that $h_{\mu}(f)>h^{s}(f)$, then $\lambda_{c}(f,\mu)\leq h^{s}(f)-h_{\mu}(f)<0$.
@@
Suppose that the central Lyapunov exponent of $\mu$ is non-negative, i.e., $\lambda^{c}(\mu,f)\geq 0$. Applying (3.2) to the stable metric entropy and considering $f^{-1}$, we obtain
$h_{\mu}(f)=h^{s}_{\mu}(f).$
Since by hypothesis $h_{\mu}(f)>h^{s}(f)$, we get $h_{\mu}^{s}(f)>h^{s}(f)$, which contradicts the variational principle for stable entropy. ∎
By $h_{\mu}(f)>h^{s}(f)$ and Claim 1, we have that $\lambda_{c}(f,\mu)<0$, and therefore, by (3.1),
@@
It is worth noting that under the hypotheses of Lemma 3.1, every ergodic measure with metric entropy bigger than the stable entropy is hyperbolic and the metric entropy coincides with the unstable metric entropy. Furthermore, the central Lyapunov exponent is negative, so its $s$-index is $\operatorname{dim}(E^{s})+1$. This leads us to inquire whether two measures with large entropy exhibit homoclinic relation as defined in the definition 2.5.
@@
Observe that the condition $P^{u}(f,\phi)>P^{s}(f,\phi)$ is not sufficient to guarantee positive entropy for the equilibrium states of $(f,\phi)$. When $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$, every measure of maximal entropy $\mu$ satisfies $P_{\mu}(f,\phi)>h^{s}(f)+\sup\phi$. So, we can state the following
@@
Let $\phi:M\to{\mathbb{R}}$ be a continuous potential such that $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$. If $\mu$ is an ergodic measure such that $P_{\mu}(f,\phi)>h^{s}(f)+\sup\phi$ then $\lambda_{c}(f,\mu)\leq h^{s}(f)-h_{\mu}(f)<0$.
@@
If $h^{u}(f)>a_{1}>a_{2}>h^{s}(f)$, then there exists a $C^{1}$ neighborhood ${{\mathcal{U}}}$ of $f$ such that for every $g\in{{\mathcal{U}}}\cap{{\rm Diff}}^{1+}(M)$ the following holds:
  1. (1)
    $h^{u}(g)>a_{1}>a_{2}>h^{s}(g)$,
  2. (2)
    the function $h:{{\rm Diff}}^{1+}(M)\cap\mathcal{U}\to\mathbb{R}$ is continuous,
  3. (3)
    if $\phi:M\to\mathbb{R}$ is a Hölder continuous potential with $a_{1}=\sup\phi$ and $a_{2}=\inf\phi$, the function $P(.,\phi):{\rm Diff}^{1+}(M)\cap\mathcal{U}\to\mathbb{R}$ is continuous.
@@
The lemma is a consequence of [30, Corollaries 3.6 and 3.7]. ∎
@@ 4. Invariant manifolds for measures with large entropy
In this section, we will investigate the measure of the set of points with invariant manifolds of uniform size from the perspective of measures with high entropy.
@@ 4.1. Invariant Manifolds with Uniform Size
We verified that ergodic measures with large entropy are hyperbolic measures, see Lemma 3.1. Almost every point of a hyperbolic measure possesses stable and unstable invariant manifolds, see Section 2.2. However, a crucial point to note is the absence of uniform control over the size of these invariant manifolds. Consequently, there may exist points that are arbitrarily close to each other such that their corresponding unstable and stable manifolds do not intersect. This suggests the potential existence of an infinite number of maximal entropy measures that are not homoclinically related. To avoid this kind of problem, in this section, we construct, for partially hyperbolic diffeomorphisms with central dimension one a set where the Pesin invariant manifolds have uniform size.
Recall that $f:M\to M$ denotes a $C^{1+}$ partially hyperbolic diffeomorphism with $1$-dimensional central direction with non trivial stable and unstable bundles, that is, ${\operatorname{dim}}(E^{\sigma})>0(\sigma=s,u)$. Given $01>\rho\geq\xi>$ 0 where $\xi$ is as in Definition 2.10, define
(4.1) $\Lambda_{\rho}(f)=\Lambda_{\rho}:=\{x\in M:\forall n\geq 1\mbox{ it holds }\|Df^{n}(x)_{|E^{c}}\|\leq\rho^{n}\}.$
@@
and the splitting $E^{cs}(x)\oplus E^{u}(x)$ varies continuously. Therefore, the classical result presented in [18, 28] guarantees the existence of invariant manifolds of uniform size. More precisely, there exists $\delta_{r}=\delta_{r}(\rho)$ and $\delta_{l}=\delta_{l}(\rho)$ such that for every $x\in\Lambda_{\rho}(f)$ unstable and stable manifold have a size larger than $\delta_{r}$ and if $x,y\in\Lambda_{\rho}(f)$ and $d(x,y)<\delta_{l}$ then $W^{u/s}_{\delta_{r}}(x)\pitchfork W^{s/u}_{\delta_{r}}(y)$ is non empty. Furthermore, the invariant manifolds exhibit continuous variations concerning both the diffeomorphisms involved and the chosen points. In other words, given $f$ and $g$ are $C^{1}$ close enough in the $C^{1}$ topology, for any $x$ in $\Lambda_{\rho}(f)$ and $y$ in $\Lambda_{\rho}(g)$ where $x$ and $y$ are close, the invariant manifolds $W^{s/u}_{\delta_{r}}(x)$ are also close to $W^{s/u}_{\delta_{r}}(y)$ within the $C^{1}$ topology.
4.2. Measuring $\Lambda_{\rho}$
Now, we know that points in $\Lambda_{\rho}(f)$ have invariant manifolds of uniform size. We will use this set to prove that there exists a finite number of homoclinic classes containing measures with large entropy. Therefore, we aim to demonstrate that the set $\Lambda_{\rho}(f)$ possesses a positive measure for measures with significant entropy. To achieve our goals, we require the following version of the Pliss Lemma [40].
@@ 4.2. Measuring $\Lambda_{\rho}$
[10, Lemma 3.1] For any $\varepsilon>0$, $\alpha_{1}<\alpha_{2}$ and any sequence $(a_{i})\in(\alpha_{1},+\infty)^{{\mathbb{N}}}$ satisfying
@@
    for any $l\geq 1$ and $n>n_{l}$, one has $\frac{a_{n_{l}}+\cdots+a_{n-1}}{n-n_{l}}\leq\alpha_{2}+\varepsilon$,
@@
We define $m(f):=\min_{x\in M}\{\log\|Df(x)_{|E^{c}_{x}}\|\}$. Note that $m(f)$ is negative because there are points with central Lyapunov exponent negative and it varies continuously with $f$ on the $C^{1}$ topology.
@@
Suppose that $h^{u}(f)>h^{s}(f)$. If $\mu\in{\mathcal{M}}_{f}^{e}$ with $h_{\mu}(f)>h^{s}(f)$ then for every $e^{\lambda_{c}(f,\mu)}<\rho<1$ we have that $\mu(\Lambda_{\rho})\geq\frac{\log(\rho)-\lambda_{c}(f,\mu)}{\log(\rho)-m(f)}.$
@@
Let $\mu\in{\mathcal{M}}_{f}$ be an ergodic measure with $h_{\mu}(f)>h^{s}(f)$. Since $h^{u}(f)>h^{s}(f)$ and $h_{\mu}(f)>h^{s}(f)$, Lemma 3.1 implies that $\lambda_{c}(f,\mu)<h^{s}(f)-h_{\mu}(f)<0$. Now, let $10<e^{\lambda_{c}(f,\mu)}<\rho<$ 1 and define the set
$\Lambda(f)=\Lambda_{\rho}\cap\mathcal{R}\cap\Lambda_{\mu},$
where $\mathcal{R}$ is the Lyapunov regular set and $\Lambda_{\mu}$ are the points in $M$ with central Lyapunov exponent equal to $\lambda_{c}(f,\mu)$. Note that $\mu(\Lambda(f))=\mu(\Lambda_{\rho})$.
@@
Consequently, we obtain $\|Df^{n}(f^{n_{l}}(x))_{|E^{c}(f^{n_{l}}(x))}\|<\rho^{n}$ for every $n\in{\mathbb{N}}$. Thus, for each $l$, $f^{n_{l}}(x)\in\Lambda(f)$. Applying Birkhoff’s theorem and the Pliss Lemma, we arrive at the following estimate:
(4.2) $\mu(\Lambda_{\rho})\geq\limsup_{l\to\infty}\frac{l}{n_{l}}\geq\frac{\log(\rho)-\lambda_{c}(f,\mu)}{\log(\rho)-m(f)}>0.$
@@
Let $\phi:M\to\mathbb{R}$ be a continuous potential such that $h^{u}(f)>\sup\phi>\inf\phi>h^{s}(f)$. For every $a$ such that $h^{u}(f)+\inf\phi>a>h^{s}(f)+\sup\phi$ and $\frac{h^{s}(f)-a+\inf\phi}{2}<\log(\rho)<1$ it holds
@@
Now, consider $\log(\rho)>\frac{h^{s}(f)+\sup\phi-a}{2}$. So, $e^{\lambda_{c}(g,\mu)}<\rho<1$, and therefore, by Corollary 4.4: